OpenStudy (anonymous):

Find derivative form of (1/(x^2)(x^2+5x+4))

OpenStudy (anonymous):

sorry messed up h/o

OpenStudy (anonymous):

\[1/(x^2-2)(x^2+5x+4)\]

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}~\frac{1}{(x^2-2)(x^2+5x+4)}\]All in the denominator?

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{x^2-2}(x^2+5x+4)\quad ?\)

OpenStudy (anonymous):

yeah thats the problem sorry for not using more parentheses. What zepdrix said...

OpenStudy (anonymous):

sidenote... is there a code list for use with the writing of equations?

zepdrix (zepdrix):

You can rewrite the expression like this if you're more comfortable using the `product rule`: \[\large\bf\sf \frac{d}{dx}~\frac{1}{(x^2-2)(x^2+5x+4)}\quad=\quad \frac{d}{dx}~(x^2-2)^{-1}(x^2+5x+4)^{-1}\]Otherwise you have to `quotient rule` and then `product rule` within your quotient rule. Both methods are gonna be a little messy. Just depends which approach you want to take.

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{(x^2-2)(x^2+5x+4)}\implies [(x^2-2)(x^2+5x+4)]^{-1}\) I'd think chain-rule it and then just use the product rule

OpenStudy (anonymous):

[4x^2+6]^-1

OpenStudy (anonymous):

there are multiple choice answers i'll post them h/o

zepdrix (zepdrix):

You can google latex to find out more information about the coding. The easiest way to get comfortable with it though.... If you see some code and want to know how someone did it, right click the code > show math as > tex commands that will show you how they input the code. Or you can use the special buttons in the equation tool to get familiar with the code words

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I still don't know the answer